OpenStudy (anonymous):

Brett has determined a function f(x) that shows the exponential growth of the number of shoes Larae owns each year. Explain how the f^−1(x) can be found and what f^−1(132) means.

OpenStudy (sidsiddhartha):

f^(-1) [x] is the inverse function of f(x), it can be found by replacing the x and y values

OpenStudy (sidsiddhartha):

take a example \[f(x)=x+3\] \[y=x+3\] then the inverse will be x = y + 3 x - 3= y \[f^{-1}(x)=x-3\] like this

OpenStudy (anonymous):

ohk so how would I find out what f^−1(132) means?

OpenStudy (sidsiddhartha):

for f^{-1}(132) you input 132 into the x value(s) of the inverse function and evaluate it. then you would plot the points on a graph where x is 132 and y is the result of 132 input in the function

OpenStudy (anonymous):

is that all I do?

OpenStudy (sidsiddhartha):

yup that'll be all

OpenStudy (anonymous):

thanks!

OpenStudy (sidsiddhartha):

yw!!